OpenStudy (jocelynstalnaker):

So I got a question wrong on my test and trying to figure out how to solve it. Help?

OpenStudy (jocelynstalnaker):

A rectangle has a length of the fifth root of 16 inches and a width of 2 to the 1 over 5 power inches. Find the area of the rectangle. 2 to the 3 over 5 power inches squared 2 to the 4 over 5 power inches squared 2 inches squared 2 to the 2 over 5 power inches squared

OpenStudy (jocelynstalnaker):

I put the first one.

OpenStudy (anonymous):

The correct answer is 2 in. squared.

OpenStudy (jocelynstalnaker):

Ok, but how did you find that out?

OpenStudy (anonymous):

Write down what you know. \[L = \sqrt[5]{16}\]\[W = 2^{\frac{ 1 }{ 5 }}\] Figure out what you are asked to solve. Area of a Rectangle \[A_{rectangle} = Length \times width\] Plug and Chug.

OpenStudy (anonymous):

ok... so you have to rewrite 16 as a power of 2 so its  use the previous rule.. then you can multiply the same base and add the powers..